OpenStudy (anonymous):

whats the slope of the line that passes through the points (-6,1) and (4,-4)

OpenStudy (anonymous):

use the M=(y2-y1)/(x2-x1) to find the slope

OpenStudy (anonymous):

slope=(y2-y1)/(x2-x1_

OpenStudy (anonymous):

(-6,1) is: x1,y1 (4,-4) is; x2 , y2

OpenStudy (anonymous):

ok so i have m=(-4-1)(4-6) what do i do now

OpenStudy (anonymous):

m=-5/-2 m=5/2

OpenStudy (anonymous):

you want to find slope that all?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Ok, you done

OpenStudy (anonymous):

but the choices are only -2, 1/2, 2, or -1/2